OpenStudy (anonymous):

According to the manual, a battery in a cellular phone loses 2% of its charge each day. Assume the battery is 100% charged. a) Write an equation to represent the percent charge, P, as a function of the number of days, d, since the battery was charged. b) Determine the number of days until the battery in only 50% charged.

OpenStudy (jamesj):

Ok, after day 1, what percentage is the battery charged? 98% What's that as a fraction?

OpenStudy (anonymous):

49/50

OpenStudy (jamesj):

Good and as a decimal it's 0.98 Now what's the charge after two days? It is *not* 1 - 0.02 -0.02 = 0.96. It is something else. What is it?

OpenStudy (jamesj):

It is 98% of what was available on day 1. I.e., 98% of 98%. That is, it is (0.98)(0.98) = (0.98)^2

OpenStudy (jamesj):

make sense?

OpenStudy (anonymous):

ya

OpenStudy (jamesj):

Hence what is the charge after 3 days?

OpenStudy (anonymous):

(0.98)^2 (0.98)

OpenStudy (jamesj):

Yes, which can be simplified to what?

OpenStudy (anonymous):

0.941192

OpenStudy (jamesj):

No, I mean it can be simplified to (0.98)^3 What is the charge therefore after n days? It is (0.98)^n To answer the last part of your question, you have to find the n such that (0.98)^n = 0.5 or the integer for which (0.98)^n is just less than 0.5 For that you will need logarithms.

OpenStudy (anonymous):

0.5^2

OpenStudy (anonymous):

=0.25